OpenStudy (anonymous):

Solve. Round results to the nearest thousandth. 2x2 + 3x - 7 = 0

OpenStudy (anonymous):

-1000

OpenStudy (campbell_st):

you need the general quadratic formula. \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 2, b = 3 and c = -7 then \[x = \frac{-3 \pm \sqrt{3^2 - 4 \times2\times(-7)}}{2\times2}\] \[x = \frac{-3 \pm \sqrt{67}}{4}\] you need to evaluate \[x = \frac{-3 + \sqrt{67}}{4}...and... x= \frac{-3 - \sqrt{67}}{4}\] correct to 3 decimal places.

OpenStudy (anonymous):

but it is not quadratic

OpenStudy (campbell_st):

\[2x^2 + 3x - 7 = 0\] seems to be a quadratic to me...

OpenStudy (anonymous):

its 2 multiplied by 2

OpenStudy (anonymous):

if it is \[2x^{2}\] it has to be 2x^2

OpenStudy (shubhamsrg):

if you're using that much common sense @izanagi.wielder ,,tell me this,,why would anyone give the ques as 2*x*2 instead of 4x ?? makes sense?

OpenStudy (anonymous):

its a book copied question

OpenStudy (anonymous):

well it will be better to the questioner

OpenStudy (anonymous):

:D